a car completes a journey at an average speed of 40 m.p.h. at what speed must it travel on the return journey if the average speed for the complete journey (out and back) is 60 m.p.h.?
Answer by Alan3354(69443)

a car completes a journey at an average speed of 40 m.p.h. at what speed must it travel on the return journey if the average speed for the complete journey (out and back) is 60 m.p.h.?
--------------
d/40 + d/r = 2d/60
1/40 + 1/r = 1/30
(r+40)/40r = 1/30
r+40 = 4r/3
3r+120 = 4r
r = 120 mi/hr